OpenStudy (anonymous):

Find x ln(7x+16) + ln 2 = 2lnx

OpenStudy (saifoo.khan):

hint: \[\ln a + \ln b = \ln(a*b)\]

OpenStudy (anonymous):

14x+32 = 2lnx ?

OpenStudy (anonymous):

\[In(14x+32)=2~In~x\]

OpenStudy (anonymous):

so 14x + 32 = x/2 ?